OpenStudy (anonymous):

help me on this one i want to understand it more (diffrent question) x − 7 = 14

OpenStudy (mathstudent55):

You need to solve the equation x - 7 = 14?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

x − 7 = 14 First you need to cancel out 7 so x stands alone. And to do that, you need to do the opposite of of the operation (-) What you do on one side, you do to the other. So: x − 7 = 14 +7 +7 ------------- x= ?

OpenStudy (anonymous):

not sure if i got it right i got 21

OpenStudy (anonymous):

it's correct ^^ You gotta believe in yourself (:

OpenStudy (mathstudent55):

Look at the left side and see what is being done to x. 7 is being subtracted from x. To undo a subtraction, you do the opposite operation which is addition. That means you need to add 7 to x - 7 to eliminate the subtraction of 7. x - 7 + 7 = x + 0 = x. What you do to one side of an equation, you must also do to the other side also. Since you need to add 7 to the left side, you must add 7 to the right side too. x - 7 = 14 +7 +7 --------------- x = 21 The solution is x = 21